how do I simplify this:
2+3/4
______
2-3/8

Answer by priya(51)

First try to simplify the numerator and denominator seperately..

we get,

(8+3)/4 - numerator
=11/4
(16 - 3)/8 - denominator.
= 13/8

therefore, (11/4)/(13/8)
= (11* 8)/(4*13)
= 22/13 is the simplified form.
want to learn maths online..visit my site!!
priya